OpenStudy (anonymous):

3+sqrt{z-6}=sqrt{z+9}

OpenStudy (anonymous):

square both sides \[9+6\sqrt{z-6}+z-6 = z+9\] \[3+z+6\sqrt{z-6}=z+9\] \[6\sqrt{z-6}=6\] \[\sqrt{z-6}=1\] Square both sides again \[z-6=1\] z=7